OpenStudy (anonymous):

How can 1/5x − 2 = 1/3x + 8 be set up as a system of equations?

OpenStudy (anonymous):

If you say that each equation is equal to some y-value, then you can get a system. I assume you mean 1/5x to be \(\frac{1}{5}x\) and 1/3x to be \(\frac{1}{3}x\), correct?

OpenStudy (anonymous):

yes youre correct

OpenStudy (anonymous):

Alright, then let's just say each side of the equality is equal to y. Then we can have this: \(y = \frac{1}{5}x -2\) \(y=\frac{1}{3}x + 8\)

OpenStudy (anonymous):

okay sounds good

OpenStudy (anonymous):

can you show me whats next

OpenStudy (anonymous):

Well, that's just the set up in order to turn it into a system of equations. Did you need to solve it as well?

OpenStudy (anonymous):

well let me show you the answer choices

OpenStudy (anonymous):

5y-5x=-10 3y-3x=24 5y-5x=-10 3y+3x=24 5y+x=-10 3y+x=24 5y-x=-10 3y-x=24

OpenStudy (anonymous):

Oh, okay, so then we have to turn it into some sort of general form. Well, if we start with \(y = \frac{1}{5}x-2\), we first want to get rid of the fraction. Do you know how to eliminate fractions in these equations?

OpenStudy (anonymous):

i knnow you would have to multiply

OpenStudy (anonymous):

Yeah. You would multiply by an LCD of any fractions in the equation. Of course the only denominator we have is a 5, so we just multiply everything by 5. Are you comfortable doing that?

OpenStudy (anonymous):

can you please do it for me. i already tried

OpenStudy (anonymous):

Well, you see why multiplying everything by 5 is what we need to do, yes?

OpenStudy (anonymous):

yes i do see that

OpenStudy (anonymous):

So if I do this: \(5*y = 5*\frac{1}{5}x - 5*2\) I get \(5y = x - 10\) That okay on your end?

OpenStudy (anonymous):

ok.... so which one of my answer choices would that be?

OpenStudy (anonymous):

Well, all of your answer choices have the 5y part as being positive and the number by itself. So that means I want to subtract x from both sides. \(5y = x - 10\) \(-x -x\) \(5y - x = -10\)

OpenStudy (anonymous):

okay thank you so much

OpenStudy (anonymous):

You're welcome